In diamminesilver(I) dicyanoargentate(I), the oxidation state of silver in the cationic complex is mentioned as a Roman numeral in the name but the oxidation state of potassium in is not mentioned as a Roman numeral in potassium trioxalatochromate(III), Why the difference?

Grade:12

1 Answers

TEJA KRISHNA
4 years ago

Talking about the oxidation state We take the average oxidation state of silver of the cationic part :[Ag(NH3)2] and anionic part :[Ag(CN)2]

Let oxidation state of Ag be X

Cationic part : X + 2(0) = 0 => X= 0 —-(1)

Anionic part : X + 2(-1) = 0 => X = 2—-(2)

Average of (1) and (2) 0+2/ 2 = 1

Therefore X= 1 .